OpenStudy (anonymous):

sqrt of 3x + 4 minus sqrt of 2x - 7 equals 3

OpenStudy (anonymous):

\[\sqrt{3x+4} - \sqrt{2x-7} = 3\]

OpenStudy (anonymous):

Could you help me with this problem? :)

OpenStudy (anonymous):

Yes. Any ideas how to start? I am a teacher and I will try to drag it out of you rather than just tell you. I won't have you suffer too much though.

OpenStudy (anonymous):

use the difference of two squares

OpenStudy (anonymous):

I figured you could square both sides first?

OpenStudy (anonymous):

Separate the roots first by adding the subtracted one to both sides then let the squaring begin.

OpenStudy (anonymous):

so add 2 x + 7

OpenStudy (anonymous):

\[\sqrt{3x+4} = 3 + \sqrt{2x-7}\]

OpenStudy (anonymous):

So how would I go about getting rid of the 3

OpenStudy (anonymous):

You have to square both sides. The right is a binomial being squared. Careful. (3 + sqrt)(3 + sqrt)

OpenStudy (anonymous):

So it would be 3x + 4 = ( 3 + sqrt of 2x - 7 ) ( 3 + sqrt of 2x - 7)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay so then the next step would be toooo...

OpenStudy (anonymous):

FOIL the right side. Then, believe it or not gather like terms so sqrt is alone and square both sides again.

OpenStudy (anonymous):

Okay I offically suck at foiling

OpenStudy (anonymous):

3*3 + 3 * sqrt + 3 * sqrt + sqrt ^ 2

OpenStudy (anonymous):

So would it be 9 + 3 sqrt of 2x - 7 + sqrt of 2x-7 squared

OpenStudy (anonymous):

\[3x +4 =9 + 6\sqrt{2x-7}+2x+7\] Simplify and isolate the sqrt

OpenStudy (anonymous):

So would I add the 3x + 4 and the 2x + 7

OpenStudy (anonymous):

Gather terms on the right. Move anything not sqrt to the left. What do you get?

OpenStudy (anonymous):

sqrt of 2x - 7 + 5 x + 11 = 22 ? that's probably wrong

OpenStudy (anonymous):

3x + 4 = 2x + 16 + sqrt

OpenStudy (anonymous):

Okay would I square both sides now?

OpenStudy (anonymous):

You there

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

Gather all but 6 sqrt on left side.

OpenStudy (anonymous):

So would I add the 3x+4 and 2x+16

OpenStudy (anonymous):

(3x+4) - (2x+16) to gather everything to the left

OpenStudy (anonymous):

So then it would be 3x + 4 - 2 x + 16 = sqrt of 2x-7 ?

OpenStudy (anonymous):

-16 and 6sqrt

OpenStudy (anonymous):

What? I'm more visual learner..

OpenStudy (anonymous):

Where did you get a 6 from???

OpenStudy (anonymous):

Way up where we did foil you see 6 * sqrt.

OpenStudy (anonymous):

So how did you get -16 and 6 sqrt

OpenStudy (anonymous):

Big picture coming. Hang on.

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

Any luck??

OpenStudy (anonymous):

I am getting there, it's just getting the visual you need takes some time.

OpenStudy (anonymous):

I had a sign error. How disgusting!

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):